Father is four times older than his son. In 14 years the father will be twice as old. What are their current ages?
Answer by ikleyn(52795)

.
In English,  "four times older" means "five times as old as".


So, if "x" is the son's current age, then the father's current age is 5x, according to the condition and correct interpretation.


In 14 years . . . 


5x + 14 = 2*(x+14)


5x + 14 = 2x + 28

5x - 2x = 28 - 14 

3x = 14  ====>  x = 14/3  is not an integer number.


Hence, the problem formulation is DEFECTIVE, since the ages in the age problems are ALWAYS (TRADITIONALLY !) integer numbers.


It means that you incorrectly use and interpret this figure of speech "four times older".



I understand it VERY WELL, since myself made this mistake, until I learned the truth from other tutors at this forum.

So, I wish you to IMPROVE YOUR ENGLISH in this part AS SOON AS YOU CAN.


It is especially important, since you try to create your own problem/problems directly translating from your native language.

---------------

There is a bunch of lessons on age word problems
    - Age problems and their solutions
    - A fresh formulation of a traditional age problem
    - Really intricate age word problem
    - Selected age word problems from the archive
    - Age problems for mental solution
in this site.

Read them and become an expert in solving age problems.

Also, become familiar with terminology and using words.

Also, you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Age word problems".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.


                    H a p p y   l e a r n i n g  ! !